OpenStudy (aurorab):

Solve: -4(x + 10) - 6 = -3(x - 2)

OpenStudy (anonymous):

Solve for x: -6-4 (x+10) = -3 (x-2) Distribute -4 over x+10. -4 (x+10) = -40-4 x: -40-4 x-6 = -3 (x-2) Group like terms in -4 x-40-6. Grouping like terms, -4 x-40-6 = -4 x+(-6-40): -4 x+(-6-40) = -3 (x-2) Evaluate -6-40. -6-40 = -46: -46-4 x = -3 (x-2) Write the linear polynomial on the left hand side in standard form. Expand out terms of the right hand side: -4 x-46 = 6-3 x Move terms with x to the left hand side. Add 3 x to both sides: 3 x-4 x-46 = (3 x-3 x)+6 Look for two terms that sum to zero. 3 x-3 x = 0: 3 x-4 x-46 = 6 Combine like terms in 3 x-4 x-46. 3 x-4 x = -x: -x-46 = 6 Isolate terms with x to the left hand side. Add 46 to both sides: (46-46)-x = 6+46 Look for two terms that sum to zero. 46-46 = 0: -x = 6+46 Evaluate 6+46. 6+46 = 52: -x = 52 Multiply both sides by a constant to simplify the equation. Multiply both sides of -x = 52 by -1: -x/(-1) = -52 Any nonzero number divided by itself is one. (-1)/(-1) = 1: Answer: | | x = -52

OpenStudy (aurorab):

thanks! and great detail! it really helped

OpenStudy (anonymous):

Np